### v1

# 22mm22mm25mm25mm

Abstract.
This article studies canonical forms derived from the finest simultaneous block diagonalization of a set of symmetric matrices via congruence. Our technique relies on Harrison’s center theory, which is extended from a single higher degree form to multiple quadratic forms, hence a set of symmetric matrices. The algebraic structures of centers and the bijective relationship between the simultaneous block diagonalization via congruence and complete sets of orthogonal idempotents of centers are investigated. We provide an algorithm that mainly uses standard linear algebra tasks and several examples to demonstrate its effectiveness. In addition, this technique can be extended verbatim to the simultaneous block diagonalization of a set of Hermitian matrices via *-congruence.
Key words and phrases:
simultaneous block diagonalization via congruence, center
2020 Mathematics Subject Classification:
15A69, 15A20, 13P05
Supported by Key Program of Natural Science Foundation of Fujian Province (Grant no. 2024J02018) and National Natural Science Foundation of China (Grant no. 12371037).
1. Introduction
1.1. Problem statement
Let $\mathbbm{k}$ be a field of characteristic $\neq 2$ and $A_{1},A_{2},\ldots,A_{m}\in\mathbbm{k}^{n\times n}$ be a set of symmetric matrices. We are interested in determining the canonical forms of the finest simultaneous block diagonalization
$P^{T}A_{i}P$ $=\begin{pmatrix}B_{i1}&&\\ &\ddots&\\ &&B_{it}\end{pmatrix}$
via congruence for all $1\leq i\leq m$, where $P\in GL_{n}(\mathbbm{k})$ is an invertible matrix, $B_{ij}\in\mathbbm{k}^{n_{j}\times n_{j}}$ and $\sum_{j=1}^{t}n_{j}=n$. This problem is called simultaneous block diagonalization via congruence (SBDC) of a set of symmetric matrices. If $t=n$, $A_{1},A_{2},\ldots,A_{m}$ are simultaneously diagonalized and it is referred to as simultaneous diagonalization via congruence (SDC). The SBDC of multiple symmetric matrices are classical problems in linear algebra and have been of interest to many researchers. By reformulating the original problem into one with less complexity, large-scale problems can be solved more efficiently. Thus, the SBDC has found applications in many areas, including quadratic programming, variational analysis and signal processing, see [3, 15, 23] and references therein.
1.2. Previous work
As a special case, the SDC is less complex and has been studied thoroughly. Early work on the SDC of two real symmetric matrices was initialized in 1868 by Weierstrass [25] and pairs of quadratic forms were developed by authors like Finsler [7] and Calabi [4]. In 2007, Hiriart-Urruty [10] posed an open question about the SDC of a set of symmetric matrices. Until recently, Jiang and Li [15] extended the SDC to more than two real symmetric matrices, which was further investigated by [19]. Some papers, such as [1, 11, 16], considered the SDC of two Hermitian matrices and it was expanded to a set of Hermitian matrices in [17]. In more recent work [3], authors tackled the SDC for a set of general complex symmetric matrices. Additionally, some researchers explored the SDC using quadratic forms associated with symmetric matrices. The SDC of two quadratic forms has been approached in [2, 9, 26] and was extended to the almost SDC for a set of quadratic forms in [24].
Compared to the SDC, which poses strict constraints on the matrices, the SBDC concerns the canonical forms of general sets of symmetric matrices and is much more difficult to solve. Uhlig [20] generalized the SDC of two real symmetric matrices to the SBDC of two real symmetric matrices based on their Jordan normal forms. His later papers [21, 22] further examined canonical pair forms of two real symmetric matrices and provided a new derivation. Until now, the SBDC of more than two symmetric matrices or Hermitian matrices has not been explored and remains a challenging problem.
1.3. Methods and results
Harrison [8] introduced centers for higher degree forms, which are highly effective for decomposing a higher degree form into a sum of forms in disjoint sets of variables [13, 14]. Recall that the center of a higher degree form, i.e. a homogeneous polynomial with degree $>2$, $f(x_{1},x_{2},\dots,x_{n})\in\mathbbm{k}[x_{1},x_{2},\dots,x_{n}]$ is defined as
$Z(f):=\{X\in\mathbbm{k}^{n\times n}\mid(H_{f}X)^{T}=H_{f}X\},$
where $H_{f}$ is the Hessian matrix of $f$. It was shown that the direct sum decomposition of a form is in bijection with the decomposition of the unit of its center algebra into orthogonal idempotents. A criterion and an algorithm mostly based on standard linear algebra tasks were provided to decompose arbitrary forms. Harrison’s center theory was extended to the simultaneous direct sum decompositions of a set of not necessarily homogeneous polynomials in [6]. The center of a set of multivariate polynomials $f_{i}(x_{1},x_{2},\dots,x_{n})\in\mathbbm{k}[x_{1},x_{2},\dots,x_{n}]$ for $1\leq i\leq m$ is defined as
[…]
It was proved that $Z(f_{1},f_{2},\dots,f_{m})$ is a special Jordan algebra and it was applied to the simultaneous direct sum decompositions of $f_{1},f_{2},\ldots,f_{m}$.
Due to the association between quadratic forms and symmetric matrices, we notice that the simultaneous direct sum decomposition of a set of quadratic forms is equivalent to the SBDC of a set of symmetric matrices. Thus, we extend Harrison’s center theory to the SBDC of a set of symmetric matrices. Centers of a set of quadratic forms defined in Equation (1.1) are reformulated to those of a set of symmetric matrices. The bijective relationship between the SBDC of a set of symmetric matrices and complete sets of orthogonal idempotents of their center is proven. A simple algorithm and several examples are provided. Finally, we generalize our approach for the SBDC of a set of symmetric matrices to the SBDC of a set of Hermitian matrices.
1.4. Organization of the paper
The remainder of this paper is organized as follows. In Section 2 we introduce and apply centers to the SBDC of a set of symmetric matrices and give a simple algorithm. Section 3 presents more examples to demonstrate the effectiveness of this technique. Section 4 extends the center theory to the SBDC of a set of Hermitian matrices.
2. The center and simultaneous block diagoanlizations
2.1. The center of a set of symmetric matrices
We begin by defining the center of a set of symmetric matrices. Let $A\in\mathbbm{k}^{n\times n}$ be a symmetric matrix and $f(x_{1},x_{2},\dots,x_{n})\in\mathbbm{k}[x_{1},x_{2},\dots,x_{n}]$ be its associated quadratic form. Since $f$ can be represented as $f(x)=x^{T}Ax$, where $x=(x_{1},x_{2},\dots,x_{n})^{T}$, it is easy to work out $H_{f}=2A$. Thus, we can reformulate the center of a set of quadratic forms in Equation (1.1) to that of a set of symmetric matrices.
Definition 2.1.
Let $A_{1},A_{2},\dots,A_{m}\in\mathbbm{k}^{n\times n}$ be a set of symmetric matrices. The center of these matrices is defined as
[…]
Remark 2.2.
Under the assumption that the characteristic of $\mathbbm{k}$ is not equal to $2$, a set of symmetric matrices and their associated quadratic forms share the same center. For clarity, we focus on symmetric matrices instead of quadratic forms in the rest of this article.
Example 2.3.
Consider symmetric matrices $A_{1}=\begin{pmatrix}-2&2&-2\\ 2&2&0\\ -2&0&-1\end{pmatrix}$ and $A_{2}=\begin{pmatrix}5&7&-1\\ 7&5&1\\ -1&1&-1\end{pmatrix}$. According to Equation (2.1), we compute the center of $A_{1}$ and $A_{2}$ as
$Z(A_{1},A_{2})=\left\{\begin{pmatrix}a-5b-c-d+5e&-a-b-5d+e&a-2b-d-e\\ 6a+6b+c-6d-6e&6a&3b\\ 2c&12d&6e\end{pmatrix}\vline\,a,b,c,d,e\in\mathbbm{k}\right\}.$
Example 2.4.
Given the following symmetric matrices
$A_{1}=\begin{pmatrix}1&1&1\\ 1&1&0\\ 1&0&1\end{pmatrix},\quad A_{2}=\begin{pmatrix}-2&2&-2\\ 2&2&0\\ -2&0&-1\end{pmatrix},\quad A_{3}=\begin{pmatrix}1&3&-2\\ 3&3&0\\ -2&0&2\end{pmatrix}.$
Their center is calculated as
$Z(A_{1},A_{2},A_{3})=\left\{\begin{pmatrix}38a+22b&0&24a\\ -19a-11b+11c&22c&-12a\\ 22a&0&22b\end{pmatrix}\vline\,a,b,c\in\mathbbm{k}\right\}.$
2.2. Jordan algebras
The center of a set of symmetric matrices in Equation (2.1) may not have an associative algebraic structure as the center of a single higher degree form. However, we observe that it has a special Jordan algebraic structure as discussed below. Some necessary definitions of Jordan algebras are included and more are given in [18].
Definition 2.5.
A Jordan algebra is a non-associative algebra, which consists of a vector space $J$ with a bilinear product $X\odot Y$ that satisfies the following identities:
$X\odot Y=Y\odot X,\quad(X^{2}\odot Y)\odot X=X^{2}\odot(Y\odot X).$
Let $A$ be an associative algebra over $\mathbbm{k}$. We attain a Jordan algebra $A^{+}$ with a symmetrized product
$X\odot Y=\frac{1}{2}(XY+YX),$
which is commutative and satisfies the above identities. A special Jordan algebra is a Jordan algebra that is isomorphic to a subalgebra of some $A^{+}$.
An idempotent is an element $e\in J$ that satisfies $e\odot e=e$. Two idempotents $e_{1}$ and $e_{2}$ are orthogonal if $e_{1}\odot e_{2}=0$. In a unital Jordan algebra $J$, a set of idempotents $e_{1},e_{2},\dots,e_{s}$ are referred to as a complete set of orthogonal idempotents if
$e_{i}\odot e_{i}=e_{i},\,\forall i,\quad e_{i}\odot e_{j}=0,\,\forall i\neq j,\quad\sum_{i=1}^{s}e_{i}=1.$
Remark 2.6.
If $e_{1},e_{2},\dots,e_{s}$ are a complete set of orthogonal idempotents in a unital Jordan subalgebra $J$ of $({\mathbbm{k}^{n\times n}})^{+}$, they must satisfy the following properties:
$e_{i}^{2}=e_{i},\,\forall i,\quad e_{i}e_{j}=0,\,\forall i\neq j,\quad\sum_{i=1}^{s}e_{i}=I_{n}.$
We will briefly explain $e_{i}e_{j}=0$ and the other two properties are obvious. Since an idempotent $e_{i}$ must satisfy $e_{i}^{2}=e_{i},$ there exists an invertible matrix $P$ such that $P^{-1}e_{i}P=\begin{pmatrix}I&0\\ 0&0\end{pmatrix}.$ Let $P^{-1}e_{j}P$ be represented as a block matrix $\begin{pmatrix}X&Y\\ Z&W\end{pmatrix}$. Since $e_{i}\odot e_{j}=\frac{1}{2}(e_{i}e_{j}+e_{j}e_{i})=0$ for any two orthogonal idempotents $e_{i},e_{j}$, we can see that $P^{-1}e_{j}P=\begin{pmatrix}0&0\\ 0&W\end{pmatrix}$ and thus $e_{i}e_{j}=0$.
Theorem 2.7.
Let $A_{1},A_{2},\ldots,A_{m}\in\mathbbm{k}^{n\times n}$ be a set of symmetric matrices and $Z(A_{1}$, $A_{2}$, $\ldots$, $A_{m})$ be their center. Then $(Z(A_{1},A_{2},\ldots,A_{m}),\odot)$ is a special Jordan algebra.
Proof.
By definition, any $X,Y\in Z(A_{1},A_{2},\ldots,A_{m})$ need to satisfy $X\odot Y=\frac{1}{2}(XY+YX)$. We have
$A_{i}(X\odot Y)=A_{i}\frac{(XY+YX)}{2}=\frac{(Y^{T}X^{T}+X^{T}Y^{T})}{2}A_{i}=(X\odot Y)^{T}A_{i}=[A_{i}(X\odot Y)]^{T},$
for $1\leq i\leq m$. This implies that $(Z(A_{1},A_{2},\ldots,A_{m}),\odot)$ is closed under symmetric product $\odot$ and hence a Jordan subalgebra of ${(\mathbbm{k}^{n\times n})}^{+}$. ∎
2.3. Simultaneous block diagoanlization of symmetric matrices
Based on the algebraic structures of centers $Z(A_{1},A_{2},\ldots,A_{m})$ discussed in Subsection 2.2, we simultaneously block diagonalize a set of symmetric matrices via congruence.
Theorem 2.8.
Suppose $A_{1},A_{2},\ldots,A_{m}\in\mathbbm{k}^{n\times n}$ are a set of symmetric matrices. Then
  • (1)
    There is a one-to-one correspondence between simultaneous block diagonalizations of $A_{1}$, $A_{2}$, $\ldots$, $A_{m}$ via congruence and complete sets of orthogonal idempotents of $Z(A_{1},A_{2},\ldots,A_{m})$.
  • (2)
    $A_{1},A_{2},\ldots,A_{m}$ are not simultaneously block diagonalizable via congruence if and only if $Z(A_{1}$, $A_{2}$, $\ldots$, $A_{m})$ has no nontrivial idempotent elements.
  • (3)
    $A_{1},A_{2},\ldots,A_{m}$ are simultaneously diagonalizable via congruence if and only if $Z(A_{1}$, $A_{2},\ldots,A_{m})$ has a complete set of $n$ orthogonal idempotents.
Proof.
A proof of item (1) is provided, which implies items (2) and (3). We begin by examining how centers of a set of symmetric matrices vary under the congruence transformation. Let $B_{1},B_{2},\ldots,B_{m}\in\mathbbm{k}^{n\times n}$ be a set of matrices. Suppose there exists an invertible matrix $P\in\operatorname{GL}_{n}(\mathbbm{k})$ such that $B_{i}=P^{T}A_{i}P$ for $1\leq i\leq m$. According to Equation (2.1), we have $(B_{i}Y)^{T}=B_{i}Y$ for any $Y\in Z(B_{1},B_{2},\ldots,B_{m})$. It is ready to see that $(B_{i}Y)^{T}=Y^{T}P^{T}A_{i}P=P^{T}A_{i}PY$, which shows that $PYP^{-1}\in Z(A_{1},A_{2},\ldots,A_{m})$. Consequently, we have $Z(B_{1},B_{2},\ldots,B_{m})=P^{-1}Z(A_{1},A_{2},\ldots,A_{m})P.$
Suppose symmetric matrices $A_{1},A_{2},\ldots,A_{m}$ are simultaneously block diagonalizable via congruence such that
$B_{i}=P^{T}A_{i}P=\begin{pmatrix}B_{i1}&&\\ &\ddots&\\ &&B_{it}\end{pmatrix}$
for all $1\leq i\leq m$. Since $B_{i}$ is a block diagonal matrix and can be written as $B_{i}=B_{i1}\oplus B_{i2}\oplus\cdots\oplus B_{it}$, we see that their centers can be represented as $Z(B_{1},B_{2},\ldots,B_{m})=Z(B_{11},B_{21},\ldots,B_{m1})\oplus Z(B_{12},B_{22},\ldots,B_{m2})\oplus\cdots\oplus Z(B_{1t},B_{2t},\ldots,B_{mt})$. Thus, there exists a correspondence between each identity matrix $I_{j}$ of $Z(B_{1j},B_{2j},\ldots,B_{mj})$ and an idempotent element $e_{j}$ of $Z(B_{1},B_{2},\ldots,B_{m})$. Note that $e_{j}$ is a block diagonal matrix that is written as $e_{j}=0\oplus\cdots\oplus I_{j}\oplus\cdots\oplus 0$. Since the center of $A_{1},A_{2},\ldots,A_{m}$ can be represented as $Z(A_{1},A_{2},\dots,A_{m})=PZ(B_{1},B_{2},\dots,B_{m})P^{-1}$, it is obvious that $Pe_{1}P^{-1},Pe_{2}P^{-1}$, $\ldots$, $Pe_{t}P^{-1}$ are a complete set of orthogonal idempotents of $Z(A_{1},A_{2},\ldots,A_{m})$.
Conversely, suppose there exists a complete set of orthogonal idempotents $\epsilon_{1},\epsilon_{2},\ldots,\epsilon_{t}$ in $Z(A_{1},A_{2},\ldots,A_{m})$. We have $\epsilon_{i}^{2}=\epsilon_{i}$ for all $i$ and $\epsilon_{i}\epsilon_{j}=0$ for all $i\neq j$ according to Remark 2.6. Since the $\epsilon_{i}$’s are diagonalizable matrices and they pairwise commute, $\epsilon_{1},\epsilon_{2},\ldots,\epsilon_{t}$ are simultaneously diagoanalizable. Therefore, there exists an invertible matrix $P\in\operatorname{GL}_{n}(\mathbbm{k})$ such that $P^{-1}\epsilon_{j}P=e_{j}$, which satisfy $\sum_{j=1}^{t}e_{j}=I_{n}$. Let $B_{i}=P^{T}A_{i}P$ for all $1\leq i\leq m$ as defined above. We now have $B_{i}e_{j}=(B_{i}e_{j})^{T}=e_{j}^{T}B_{i}^{T}=e_{j}B_{i}$ for $1\leq i\leq m$ and $1\leq j\leq t$. We further deduce that the $B_{i}$’s are block diagonal matrices and can be written as
$B_{i}$ $=\begin{pmatrix}B_{i1}&&\\ &\ddots&\\ &&B_{it}\end{pmatrix}$
for all $1\leq i\leq m$. Therefore, a simultaneous block diagonalization of $A_{1},A_{2},\ldots,A_{m}$ via congruence corresponds to the complete set of orthogonal idempotents $\epsilon_{1},\epsilon_{2},\ldots,\epsilon_{t}$. ∎
2.4. Algorithm for simultaneous block diagonalization
We describe the procedures of the SBDC of a set of symmetric matrices in Algorithm 1. It is a recursive process that begins with computing the center of these symmetric matrices according to Definition 2.1. Using the base matrices of the center, a pair of orthogonal idempotents $\epsilon_{1}$ and $\epsilon_{2}$ is calculated. If no orthogonal idempotent is found, the symmetric matrices are not simultaneously block diagonalizable via congruence. Otherwise, an invertible matrix $P$ is computed according to Theorem 2.8 and these symmetric matrices are simultaneously block diagonalized via congruence as a set of block diagonal matrices, each with two diagonal block submatrices.
Algorithm 1 Simultaneous block diagonalization of symmetric matrices
1:Input: a set of symmetric matrices $A_{1},A_{2},\ldots,A_{m}$
2:Output: two sets of symmetric matrices $B_{11},B_{21},\ldots,B_{m1}$ and $B_{12},B_{22},\ldots,B_{m2}$
3:Caculate the center $Z(A_{1},A_{2},\ldots,A_{m})$ by solving system $(A_{i}X)^{T}=A_{i}X$ for $1\leq i\leq m$. Select base matrices $X_{j}$ for $1\leq j\leq\operatorname{dim}{Z(A_{1},A_{2},\ldots,A_{m})}$.
4:Compute a pair of orthogonal idempotents $\epsilon_{1}$ and $\epsilon_{2}$ using $Z(A_{1},A_{2},\ldots,A_{m})$. If they do not exist, $A_{1},A_{2},\ldots,A_{m}$ are not simultaneously block diagonalizable via congruence. Otherwise, continue.
5:Compute invertible matrix $P$ such that $P^{-1}\epsilon_{1}P=\begin{pmatrix}I_{1}&\\ &0\end{pmatrix}$ and $P^{-1}\epsilon_{2}P=\begin{pmatrix}0&\\ &I_{2}\end{pmatrix}$. Then simultaneously block diagonalize $A_{1},A_{2},\ldots,A_{m}$ such that $P^{T}A_{i}P$ takes the form of a block diagonal matrix $\begin{pmatrix}B_{i1}&\\ &B_{i2}\end{pmatrix}$ for $1\leq i\leq m$. Go to step 1 and continue to simultaneously block diagonalize $B_{11},B_{21},\ldots,B_{m1}$ and $B_{12},B_{22},\ldots,B_{m2}$ via congruence separately.
Since these submatrices may be simultaneously block diagonalizable via congruence, Algorithm 1 is called again using the submatrices as input. This process occurs recursively until all output matrices are not simultaneously block diagonalizable via congruence. The resulting output matrices can then be reassembled as block diagonal matrices corresponding to the original symmetric matrices $A_{1},A_{2},\ldots,A_{m}$. During this process, the maximum number of blocks for the SBDC is also determined.
This algorithm contains several components, including computations of centers and orthogonal idempotents, which use standard tasks in linear algebra [14]. Calculating the centers of symmetric matrices requires solving a system of linear equations, which has polynomial complexity [5]. The orthogonal idempotents are obtained by imposing constraints on the rank and trace of matrix $X$ in Equation (2.1), which leads to a system of nonlinear equations. For the SDC of a set of symmetric matrices, this system only contains sparse quadratic equations that are solved efficiently in numerical experiments. We refer to [5] for more information. The complexity analysis of the SBDC of a set of symmetric matrices is still unavailable and will be addressed in future studies.
3. More examples
We provide more examples to demonstrate the procedure of this technique. The SDC of two and three symmetric matrices are shown in Examples 3.1 and 3.2, respectively. The SBDC of two and three symmetric matrices are given in Examples 3.3 and 3.4, respectively.
Example 3.1.
(Example 2.3 continued) Consider symmetric matrices
$A_{1}=\begin{pmatrix}-2&2&-2\\ 2&2&0\\ -2&0&-1\end{pmatrix},\quad A_{2}=\begin{pmatrix}5&7&-1\\ 7&5&1\\ -1&1&-1\end{pmatrix}.$
Using the center calculated in Example 2.3, we obtain base matrices of $Z(A_{1},A_{2})$ as
$X_{1}=\begin{pmatrix}1&-1&1\\ 6&6&0\\ 0&0&0\end{pmatrix},\quad X_{2}=\begin{pmatrix}-5&-1&-2\\ 6&0&3\\ 0&0&0\end{pmatrix},\quad X_{3}=\begin{pmatrix}-1&0&0\\ 1&0&0\\ 2&0&0\end{pmatrix},$
$X_{4}=\begin{pmatrix}-1&-5&-1\\ -6&0&0\\ 0&12&0\end{pmatrix},\quad X_{5}=\begin{pmatrix}5&1&-2\\ -6&0&0\\ 0&0&6\end{pmatrix}.$
Since $-X_{3}$ is a primitive idempotent, we set $\epsilon_{1}=-X_{3}$ and $\epsilon_{2}=I-\epsilon_{1}$. A pair of orthogonal idempotents is produced as
$\epsilon_{1}=\begin{pmatrix}1&0&0\\ -1&0&0\\ -2&0&0\end{pmatrix},\quad\epsilon_{2}=\begin{pmatrix}0&0&0\\ 1&1&0\\ 2&0&1\end{pmatrix}$
and we calculate an invertible matrix $P_{1}=\begin{pmatrix}-1&0&0\\ 1&1&0\\ 2&0&1\end{pmatrix}$ according to Theorem 2.8 such that
$P_{1}^{-1}\epsilon_{1}P_{1}=\begin{pmatrix}1&&\\ &0&\\ &&0\end{pmatrix},\quad P_{1}^{-1}\epsilon_{2}P_{1}=\begin{pmatrix}0&&\\ &1&\\ &&1\end{pmatrix}.$
$A_{1}$ and $A_{2}$ can then be simultaneously block diagonalized via congruence as
$P_{1}^{T}A_{1}P_{1}=\begin{pmatrix}0&&\\ &2&0\\ &0&-1\end{pmatrix},\quad P_{1}^{T}A_{2}P_{1}=\begin{pmatrix}0&&\\ &5&1\\ &1&-1\end{pmatrix},$
respectively. Since neither matrices are fully diagonalized, we continue to simultaneously diagonalize their bottom-right submatrices, which are denoted as $B_{1}=\begin{pmatrix}2&0\\ 0&-1\end{pmatrix}$ and $B_{2}=\begin{pmatrix}5&1\\ 1&-1\end{pmatrix}$. Their center is calculated as
$Z(B_{1},B_{2})=\left\{\begin{pmatrix}-3a+b&-a\\ 2a&b\end{pmatrix}\vline\,a,b\in\mathbbm{k}\right\}.$
We obtain a new pair of idempotent elements using $Z(B_{1},B_{2})$ as
$\epsilon_{3}=\begin{pmatrix}-1&-1\\ 2&2\end{pmatrix},\quad\epsilon_{4}=\begin{pmatrix}2&1\\ -2&-1\end{pmatrix}$
and derive $P_{2}=\begin{pmatrix}-1&-1\\ 2&1\end{pmatrix}$ such that $P_{2}^{-1}\epsilon_{3}P_{2}=\begin{pmatrix}1&0\\ 0&0\end{pmatrix}$ and $P_{2}^{-1}\epsilon_{4}P_{2}=\begin{pmatrix}0&0\\ 0&1\end{pmatrix}$. Thus, $B_{1}$ and $B_{2}$ are simultaneously diagonalized via congruence as
$P_{2}^{T}B_{1}P_{2}=\begin{pmatrix}-2&0\\ 0&1\end{pmatrix},\quad P_{2}^{T}B_{2}P_{2}=\begin{pmatrix}-3&0\\ 0&2\end{pmatrix},$
respectively.
Finally, we can see that $A_{1}$ and $A_{2}$ are simultaneously diagonalized via congruence as
$P^{T}A_{1}P=\begin{pmatrix}0&&\\ &-2&\\ &&1\end{pmatrix},\quad P^{T}A_{2}P=\begin{pmatrix}0&&\\ &-3&\\ &&2\end{pmatrix},$
respectively, where
$P=P_{1}\begin{pmatrix}1&0\\ 0&P_{2}\end{pmatrix}=\begin{pmatrix}-1&0&0\\ 1&-1&-1\\ 2&2&1\end{pmatrix}.$
Note that $1,1$-th entries of both transformed matrices are zero. This implies that $A_{1}$ and $A_{2}$ are degenerate matrices.
Example 3.2.
Consider the following symmetric matrices
$A_{1}=\begin{pmatrix}1&1&1\\ 1&3&3\\ 1&3&15\end{pmatrix},\quad A_{2}=\begin{pmatrix}1&1&1\\ 1&4&4\\ 1&4&20\end{pmatrix},\quad A_{3}=\begin{pmatrix}1&1&1\\ 1&5&5\\ 1&5&25\end{pmatrix}.$
We compute their center by Equation (2.1) as
$Z(A_{1},A_{2},A_{3})=\left\{\begin{pmatrix}a+b+c&a&a\\ 0&b+c&b\\ 0&0&c\end{pmatrix}\vline\,a,b,c\in\mathbbm{k}\right\}$
and its base matrices are
$X_{1}=\begin{pmatrix}1&1&1\\ 0&0&0\\ 0&0&0\end{pmatrix},\quad X_{2}=\begin{pmatrix}1&0&0\\ 0&1&1\\ 0&0&0\end{pmatrix},\quad X_{3}=\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&1\end{pmatrix}.$
A pair of orthogonal idempotents is calculated as
$\epsilon_{1}=\begin{pmatrix}1&1&1\\ 0&0&0\\ 0&0&0\end{pmatrix},\quad\epsilon_{2}=\begin{pmatrix}0&-1&-1\\ 0&1&0\\ 0&0&1\end{pmatrix}$
and we have
$P_{1}^{-1}\epsilon_{1}P_{1}=\begin{pmatrix}1&&\\ &0&\\ &&0\end{pmatrix},\quad P_{1}^{-1}\epsilon_{2}P_{1}=\begin{pmatrix}0&&\\ &1&\\ &&1\end{pmatrix},$
where $P_{1}=\begin{pmatrix}1&-1&-1\\ 0&1&0\\ 0&0&1\end{pmatrix}$. Thus, we can simultaneously block diagonalize the symmetric matrices via congruence as
$P_{1}^{T}A_{1}P_{1}=\begin{pmatrix}1&&\\ &2&2\\ &2&14\end{pmatrix},\>\quad P_{1}^{T}A_{2}P_{1}=\begin{pmatrix}1&&\\ &3&3\\ &3&19\end{pmatrix},\>\quad P_{1}^{T}A_{3}P_{1}=\begin{pmatrix}1&&\\ &4&4\\ &4&24\end{pmatrix}.$
We repeat this process using their submatrices, which are denoted as
$B_{1}=\begin{pmatrix}2&2\\ 2&14\end{pmatrix},\>\quad B_{2}=\begin{pmatrix}3&3\\ 3&19\end{pmatrix},\>\quad B_{3}=\begin{pmatrix}4&4\\ 4&24\end{pmatrix}.$
Their center is calculated as
$Z(B_{1},B_{2},B_{3})=\left\{\begin{pmatrix}a+b&a\\ 0&b\\ \end{pmatrix}\vline\,a,b\in\mathbbm{k}\right\}.$
We obtain a new pair of orthogonal idempotent $\epsilon_{3}=\begin{pmatrix}1&1\\ 0&0\end{pmatrix}$ and $\epsilon_{4}=\begin{pmatrix}0&-1\\ 0&1\end{pmatrix}$, then derive a corresponding invertible matrix $P_{2}=\begin{pmatrix}1&-1\\ 0&1\end{pmatrix}$ such that $P_{2}^{T}\epsilon_{3}P_{2}=\begin{pmatrix}1&0\\ 0&0\end{pmatrix}$ and $P_{2}^{T}\epsilon_{4}P_{2}=\begin{pmatrix}0&0\\ 0&1\end{pmatrix}$. Thus, $B_{1},B_{2},B_{3}$ are simultaneously diagonalized via congruence as
$P_{2}^{T}B_{1}P_{2}=\begin{pmatrix}2&0\\ 0&12\end{pmatrix},\>\quad P_{2}^{T}B_{2}P_{2}=\begin{pmatrix}3&0\\ 0&16\end{pmatrix},\>\quad P_{2}^{T}B_{3}P_{2}=\begin{pmatrix}4&0\\ 0&20\end{pmatrix},$
respectively. Finally, $A_{1},A_{2},A_{3}$ are simultaneously diagonalized via congruence as
$P^{T}A_{1}P=\begin{pmatrix}1&&\\ &2&\\ &&12\end{pmatrix},\>\quad P^{T}A_{2}P=\begin{pmatrix}1&&\\ &3&\\ &&16\end{pmatrix},\>\quad P^{T}A_{3}P=\begin{pmatrix}1&&\\ &4&\\ &&20\end{pmatrix},$
respectively, where
$P=P_{1}\begin{pmatrix}1&0\\ 0&P_{2}\end{pmatrix}=\begin{pmatrix}1&0&0\\ 0&2&0\\ 0&2&12\end{pmatrix}.$
Example 3.3.
Let symmetric matrices be
$A_{1}=\begin{pmatrix}1&1&0\\ 1&3&6\\ 0&6&16\end{pmatrix},\quad A_{2}=\begin{pmatrix}1&1&0\\ 1&4&8\\ 0&8&20\end{pmatrix}.$
We calculate their center according to Equation (2.1) as
$Z(A_{1},A_{2})=\left\{\begin{pmatrix}a+2b-4c+18d&a&4c-51d\\ -2b&2b-4c+18d&4b-4c+51d\\ b&c&d\end{pmatrix}\vline\,a,b,c,d\in\mathbbm{k}\right\}$
and work out a set of base matrices
$X_{1}=\begin{pmatrix}1&1&0\\ 0&0&0\\ 0&0&0\end{pmatrix},\enspace X_{2}=\begin{pmatrix}2&0&0\\ -2&2&4\\ 1&0&0\end{pmatrix},\enspace X_{3}=\begin{pmatrix}-4&0&4\\ 0&-4&-4\\ 0&1&0\end{pmatrix},\enspace X_{4}=\begin{pmatrix}18&0&-51\\ 0&18&51\\ 0&0&1\end{pmatrix}.$
Based on the algebraic structure of $Z(A_{1},A_{2})$, we compute orthogonal idempotents
$\epsilon_{1}=\begin{pmatrix}1&1&0\\ 0&0&0\\ 0&0&0\end{pmatrix},\quad\epsilon_{2}=\begin{pmatrix}0&-1&0\\ 0&1&0\\ 0&0&1\end{pmatrix}$
and
$P^{-1}\epsilon_{1}P=\begin{pmatrix}1&&\\ &0&\\ &&0\end{pmatrix},\quad P^{-1}\epsilon_{2}P=\begin{pmatrix}0&&\\ &1&\\ &&1\end{pmatrix},$
where $P=\begin{pmatrix}1&-1&0\\ 0&1&0\\ 0&0&1\end{pmatrix}$.
By simultaneously block diagonalizing $A_{1}$ and $A_{2}$ via congruence using $P$, we obtain
$P^{T}A_{1}P=\begin{pmatrix}1&&\\ &3&8\\ &8&20\end{pmatrix},\quad P^{T}A_{2}P=\begin{pmatrix}1&&\\ &2&6\\ &6&16\end{pmatrix},$
which may be further simultaneously diagonalized via congruence. We denote their bottom-right submatrices as
$B_{1}=\begin{pmatrix}3&8\\ 8&20\end{pmatrix},\quad B_{2}=\begin{pmatrix}2&6\\ 6&16\end{pmatrix},$
respectively, and compute their center as
$Z(B_{1},B_{2})=\left\{\begin{pmatrix}a&-4b\\ b&a+4b\end{pmatrix}\vline\,a,b\in\mathbbm{k}\right\}.$
Since $Z(B_{1},B_{2})\cong\mathbbm{k}[\varepsilon]/(\varepsilon^{2})$ does not have a nontrivial idempotent, $B_{1},B_{2}$ are not simultaneously diagonalizable via congruence. Thus, $A_{1},A_{2}$ are simultaneously block diagonalizable via congruence.
Example 3.4.
Suppose we have symmetric matrices
$A_{1}=\begin{pmatrix}1&2&3\\ 2&8&16\\ 3&16&33\end{pmatrix},\quad A_{2}=\begin{pmatrix}1&2&3\\ 2&6&12\\ 3&12&25\end{pmatrix},\quad A_{3}=\begin{pmatrix}1&2&3\\ 2&7&16\\ 3&16&37\end{pmatrix}.$
We compute their center as
$Z(A_{1},A_{2},A_{3})=\left\{\begin{pmatrix}a-8b+c&2a-b&3a\\ 0&-12b+c&-12b\\ 0&3b&c\end{pmatrix}\vline\,a,b,c\in\mathbbm{k}\right\}$
with base matrices
$X_{1}=\begin{pmatrix}1&2&3\\ 0&0&0\\ 0&0&0\end{pmatrix},\enspace X_{2}=\begin{pmatrix}-8&-1&0\\ 0&-12&-12\\ 0&3&0\end{pmatrix},\enspace X_{3}=\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&1\end{pmatrix}.$
Since $X_{1}$ is a primitive idempotent, we set $\epsilon_{1}=X_{1}$ and obtain a pair of orthogonal idempotents
$\epsilon_{1}=\begin{pmatrix}1&2&3\\ 0&0&0\\ 0&0&0\end{pmatrix},\quad\epsilon_{2}=\begin{pmatrix}0&-2&-3\\ 0&1&0\\ 0&0&1\end{pmatrix}.$
Thus, we compute an invertible matrix $P=\begin{pmatrix}1&-2&-3\\ 0&1&0\\ 0&0&1\end{pmatrix}$ such that
$P^{-1}\epsilon_{1}P=\begin{pmatrix}1&&\\ &0&\\ &&0\end{pmatrix},\quad P^{-1}\epsilon_{2}P=\begin{pmatrix}0&&\\ &1&\\ &&1\end{pmatrix}.$
Then $A_{1},A_{2},A_{3}$ are simultaneously block diagonalized via congruence as
$P^{T}A_{1}P=\begin{pmatrix}1&&\\ &4&10\\ &10&24\end{pmatrix},\quad P^{T}A_{2}P=\begin{pmatrix}1&&\\ &2&6\\ &6&16\end{pmatrix},\quad P^{T}A_{3}P=\begin{pmatrix}1&&\\ &3&10\\ &10&28\end{pmatrix},$
respectively. We continue to simultaneously diagonalize their submatrices via congruence, which are represented as
$B_{1}=\begin{pmatrix}4&10\\ 10&24\end{pmatrix},\quad B_{2}=\begin{pmatrix}2&6\\ 6&16\end{pmatrix},\quad B_{2}=\begin{pmatrix}3&10\\ 10&28\end{pmatrix}.$
Their center is calculated as
$Z(B_{1},B_{2},B_{3})=\left\{\begin{pmatrix}4a+b&4a\\ -a&b\end{pmatrix}\vline\,a,b\in\mathbbm{k}\right\}$
and we obtain its base matrices as
$X_{4}=\begin{pmatrix}4&4\\ -1&0\end{pmatrix},\quad X_{5}=\begin{pmatrix}1&0\\ 0&1\end{pmatrix}.$
Since a nontrivial idempotent does not exist in $Z(B_{1},B_{2},B_{3})\cong\mathbbm{k}[\varepsilon]/(\varepsilon^{2})$, $B_{1},B_{2},B_{3}$ are not simultaneously diagonalizable via congruence and $A_{1},A_{2},A_{3}$ are simultaneously block diagonalizable via congruence as shown above.
4. Simultaneous block diagonalization of Hermitian matrices
In this section, we extend our results on the SBDC of a set of symmetric matrices to the SBDC of a set of Hermitian matrices. For the convenience of readers, we recall the definition of Hermitian matrices, see [12] for more details.
Definition 4.1.
A matrix $A\in\mathbbm{C}^{n\times n}$ is Hermitian if $A^{*}=A$, where $A^{*}$ denotes the conjugate transpose of $A$.
Given a set of Hermitian matrices $A_{1},A_{2},\dots,A_{m}\in\mathbbm{C}^{n\times n}$, we are considering the problem of finding the finest simultaneous block diagonalization via *-congruence
$P^{*}A_{i}P=\begin{pmatrix}B_{i1}&&\\ &\ddots&\\ &&B_{it}\end{pmatrix}$
for all $1\leq i\leq m$, where $P\in GL_{n}(\mathbbm{C})$ is an invertible matrix, $B_{ij}\in\mathbbm{C}^{n_{j}\times n_{j}}$ and $\sum_{j=1}^{t}n_{j}=n$. Note that the center of a set of symmetric matrices defined in Equation (2.1) can be naturally extended to that of a set of Hermitian matrices.
Definition 4.2.
Let $A_{1},A_{2},\dots,A_{m}\in\mathbbm{C}^{n\times n}$ be a set of Hermitian matrices. The center of these matrices is defined as
[…]
Due to their similar algebraic structures, we may apply center $Z(A_{1},A_{2}$, $\ldots,A_{m})$ in Equation (4.1) to simultaneously block diagonalize a set of Hermitian matrices via *-congruence.
Theorem 4.3.
Suppose $A_{1},A_{2},\ldots,A_{m}\in\mathbbm{C}^{n\times n}$ are a set of Hermitian matrices and $Z(A_{1},A_{2}$, $\ldots,A_{m})$ is the center of these matrices. Then
  • (1)
    $(Z(A_{1},A_{2},\ldots,A_{m}),\odot)$ is a real special Jordan algebra.
  • (2)
    There is a one-to-one correspondence between simultaneous block diagonalizations of $A_{1},A_{2},\ldots,A_{m}$ via *-congruence and complete sets of orthogonal idempotents of $Z(A_{1},A_{2},\ldots,A_{m})$.
  • (3)
    $A_{1},A_{2},\ldots,A_{m}$ are not simultaneously block diagonalizable via *-congruence if and only if $Z(A_{1},A_{2},\ldots,A_{m})$ has no nontrivial idempotent elements.
  • (4)
    $A_{1},A_{2},\ldots,A_{m}$ are simultaneously diagonalizable via *-congruence if and only if $Z(A_{1}$, $A_{2},\ldots,A_{m})$ has a complete set of $n$ orthogonal idempotents.
Proof.
The center of a set of Hermitian matrices in Definition 4.2 uses a conjugate transpose instead of the transpose used by the center of a set of symmetric matrices in Definition 2.1. Since both of them obey the reverse-order law, one can easily derive a proof for item (1) based on the proof in Theorems 2.7. Note that $X\in Z(A_{1},A_{2},\ldots,A_{m})$, but $iX$ may not $\in Z(A_{1},A_{2},\ldots,A_{m}),$ so $Z(A_{1},A_{2},\ldots,A_{m})$ is only a real subspace of $\mathbbm{C}^{n\times n}$. Thus, $(Z(A_{1},A_{2},\ldots,A_{m}),\odot)$ is a Jordan subalgebra of ${(\mathbbm{C}^{n\times n})}^{+}$ both viewed over the reals. Similarly, proof for items (2) to (4) can be reproduced based on the proof for Theorems 2.8. ∎
The SBDC of a set of Hermitian matrices follows the same procedure as the SBDC of a set of symmetric matrices. Thus, we adopt Algorithm 1 for the SBDC of a set of Hermitian matrices and provide two examples to demonstrate the process. Example 4.4 shows the SDC of three Hermitian matrices and Example 4.5 illustrates the SBDC of two Hermitian matrices.
Example 4.4.
Consider the following Hermitian matrices
$A_{1}=\begin{pmatrix}1&1+i\\ 1-i&1\end{pmatrix},\quad A_{2}=\begin{pmatrix}2&2+2i\\ 2-2i&7\end{pmatrix},\quad A_{3}=\begin{pmatrix}-2&-2-2i\\ -2+2i&1\end{pmatrix}.$
According to Equation (4.1), their center is calculated as
$Z(A_{1},A_{2},A_{3})=\left\{\begin{pmatrix}a+b&a+bi\\ 0&b\end{pmatrix}\vline\,a,b\in\mathbb{R}\right\}.$
We then compute a pair of orthogonal idempotents based on $Z(A_{1},A_{2},A_{3})$ as
$\epsilon_{1}=\begin{pmatrix}0&-1-i\\ 0&1\end{pmatrix},\quad\epsilon_{2}=\begin{pmatrix}1&1+i\\ 0&0\end{pmatrix}.$
According to Theorem 4.3, there exists an invertible matrix $P$ such that
$P^{-1}\epsilon_{1}P=\begin{pmatrix}1&0\\ 0&0\end{pmatrix},\quad P^{-1}\epsilon_{2}P=\begin{pmatrix}0&0\\ 0&1\end{pmatrix},$
where
$P=\begin{pmatrix}1+i&1\\ -1&0\end{pmatrix}.$
Thus, we can simultaneously diagonalize these Hermitian matrices via *-congruence as
$P^{*}A_{1}P$ $=\begin{pmatrix}1+i&1\\ -1&0\end{pmatrix}\begin{pmatrix}1&1+i\\ 1-i&1\end{pmatrix}\begin{pmatrix}1+i&1\\ -1&0\end{pmatrix}=\begin{pmatrix}-1&0\\ 0&1\end{pmatrix},$
$P^{*}A_{2}P$ $=\begin{pmatrix}1+i&1\\ -1&0\end{pmatrix}\begin{pmatrix}2&2+2i\\ 2-2i&7\end{pmatrix}\begin{pmatrix}1+i&1\\ -1&0\end{pmatrix}=\begin{pmatrix}3&0\\ 0&2\end{pmatrix},$
$P^{*}A_{3}P$ $=\begin{pmatrix}1+i&1\\ -1&0\end{pmatrix}\begin{pmatrix}-2&-2-2i\\ -2+2i&1\end{pmatrix}\begin{pmatrix}1+i&1\\ -1&0\end{pmatrix}=\begin{pmatrix}5&0\\ 0&-2\end{pmatrix}.$
Example 4.5.
Given two Hermitian matrices
$A_{1}=\begin{pmatrix}0&-2-2i&-3-2i\\ -2+2i&-2&4+6i\\ -3+2i&4-6i&11\end{pmatrix},\quad A_{2}=\begin{pmatrix}-2&-2-5i&-3-i\\ -2+5i&-3&11+6i\\ -3+i&11-6i&19\end{pmatrix}.$
We can compute their center as
$Z(A_{1},A_{2})=\left\{\begin{pmatrix}a+8b&6bi&-6bi\\ 3bi&a-b&3b\\ -2bi&2b&a\end{pmatrix}\vline\,a,b\in\mathbb{R}\right\}$
and then obtain a pair of orthogonal idempotents
$\epsilon_{1}=\begin{pmatrix}6&6i&-6i\\ 3i&-3&3\\ -2i&2&-2\end{pmatrix},\quad\epsilon_{2}=\begin{pmatrix}-5&-6i&6i\\ -3i&4&-3\\ 2i&-2&3\end{pmatrix}.$
An invertible matrix $P_{1}$ is calculated such that
$P_{1}^{-1}\epsilon_{1}P_{1}=\begin{pmatrix}1&&\\ &0&\\ &&0\end{pmatrix},\quad P_{1}^{-1}\epsilon_{2}P_{1}=\begin{pmatrix}0&&\\ &1&\\ &&1\end{pmatrix},$
where
$P_{1}=\begin{pmatrix}6&-3i&2i\\ 3i&2&-1\\ -2i&-1&1\end{pmatrix}.$
Now simultaneously block diagonalize $A_{1}$ and $A_{2}$ via *-congruence as
$P_{1}^{*}A_{1}P_{1}=\begin{pmatrix}2&&\\ &-1&1+i\\ &1-i&1\end{pmatrix},\quad P_{1}^{*}A_{2}P_{1}=\begin{pmatrix}1&&\\ &-1&2+i\\ &2-i&2\end{pmatrix},$
respectively.
Next, see if the transformed matrices may be further simultaneously diagonalized via *-congruence. We denote their submatrices as
$B_{1}=\begin{pmatrix}-1&1+i\\ 1-i&1\end{pmatrix},\quad B_{2}=\begin{pmatrix}-1&2+i\\ 2-i&2\end{pmatrix},$
respectively. The center of the submatrices is calculated as
$Z(B_{1},B_{2})=\left\{\begin{pmatrix}a+b-bi&-bi\\ -b&a+bi\end{pmatrix}\vline\,a,b\in\mathbb{R}\right\}.$
By direct computation, the existence of a nontrivial orthogonal idempotent in $Z(B_{1},B_{2})$ requires
$a=\dfrac{\sqrt{3}i+3}{6},b=-\dfrac{\sqrt{3}i}{3}\quad\text{or}\quad a=\dfrac{-\sqrt{3}i+3}{6},b=\dfrac{\sqrt{3}i}{3}.$
Since such $a,b\notin\mathbb{R}$, a nontrivial idempotent does not exist. Indeed, this can also be seen by the fact that $Z(B_{1},B_{2})\cong\mathbbm{C}$ as algebras. Thus, $B_{1},B_{2}$ are not simultaneously diagonalizable via *-congruence and $A_{1},A_{2}$ are simultaneously block diagonalizable via *-congruence.